# A note on the triple product property for finite groups with abelian normal subgroups of prime index

### v5

@@ 3. Technical Results
(2) Assume $U_{0}=U\cap H\trianglelefteq G$ and let $xH,yH\in\overline{S}\cap\overline{T}$ be such that $xh,yH\neq H$ and $xH\neq yH$. Suppose there are elements $s_{x}\in S_{x}$, $s_{y}\in S_{y}$, $t_{x}\in T_{x}$, $t_{y}\in T_{y}$, $u_{0},\widetilde{u}_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}t_{y}\widetilde{u}_{0}$, or, $st_{y}\bar{u}_{0}t_{x}^{-1}=1$, where $s=s_{x}s_{y}^{-1}$ and $\bar{u}_{0}=\widetilde{u}_{0}u_{0}^{-1}$. By the normality of $U_{0}$ in $G$ there is a $\hat{u}_{0}\in U_{0}$ such that $\bar{u}_{0}t_{x}^{-1}=t_{x}^{-1}\hat{u}_{0}$, which implies that $st_{y}\bar{u}_{0}t_{x}^{-1}=st_{y}t_{x}^{-1}\hat{u}_{0}=1$. However, the TPP for $S,T,U$ implies that $t_{x}=t_{y}$, a contradiction. Thus $S_{x}^{-1}T_{x}U_{0}$, $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$, which proves part (2). ∎
4. The Main Result
@@
If $G$ is a group with an abelian normal subgroup of prime index $p$ then $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$.
@@
Let $G$ and $p$ be as given. Let $H\unlhd G$ be abelian of index $p$. Then $G/H$ is a cyclic group of order $p=|G:H|=\frac{|G|}{|H|}$ whose elements are $p$ (left) cosets $x_{0}H=H,x_{1}H,\ldots,x_{p-1}H$ of $H$ in $G$ with representatives $x_{0}=1,x_{1},\ldots,x_{p-1}\in G$ respectively. Let $(S,T,U)$ be a subgroup TPP triple of $G$ with $H$-supports $\overline{S},\overline{T},\overline{U}$ as defined in Observation 3.1. By part (2) of that Observation $\overline{S},\overline{T},\overline{U}\leq G/H$, and if $\sigma=|\overline{S}|$, $\tau=|\overline{T}|$, $\upsilon=|\overline{U}|$ then $\sigma,\tau,\upsilon\in\{1,p\}$. The sets $S_{x},T_{y},U_{z}$, as defined in Lemma 3.3 will also be used at certain points: these are defined as $S_{x}:=S\cap xH$, $T_{y}:=T\cap yH$, $U_{z}:=U\cap zH$ for any $xH,yH,zH\in G/H$. By definition, if $xH\in G/H$ then $xH\in\overline{S}\iff S_{x}\neq\emptyset$, and analagously for $T_{y}$ and $U_{z}$.
@@
Let $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$. Then $(S_{0},T_{0},U_{0})$ is a subgroup TPP triple of $H$ (by Observation 2.4), and $S_{0}T_{0}U_{0}\mathrel{\unlhd}H$ (as $S_{0}T_{0}U_{0}\leq H$ and $H$ is abelian) and $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{\sigma\tau\upsilon}\leq|H|$ and $|S||T||U|\leq\frac{\sigma\tau\upsilon}{p}|G|$ (by Proposition 2.5 and Observation 3.2).
Suppose $|S||T||U|=\sigma\tau\upsilon|S_{0}||T_{0}||U_{0}|>|G|=p|H|$. Then at least two of $\sigma,\tau,\upsilon$ must exceed $1$, that is, must be equal to $p$. Without loss of generality let $\sigma=\tau=p$, which means that $S$ and $T$ are supported on all $p$ cosets of $H$, or, equivalenty, $\overline{S}=\overline{T}=G/H$ and all sets $S_{x}$ and $T_{y}$ as defined above are non-empty. Suppose $\upsilon=1$, that is, $U=U_{0}$. Then $|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{2}}>\frac{|H|}{p}$, i.e. $\frac{|H|}{|S_{0}||T_{0}||U_{0}|}=|H:S_{0}T_{0}U_{0}|<p$, which implies that $S_{0}T_{0}U_{0}=H$ and $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=|H|$, and, using Lemma 3.3, elements $s_{x}\in S_{x}\neq S_{0}$, $t_{x}\in T_{x}\neq T_{0}$ can be chosen so that the translate $s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})=S_{0}s_{x}^{-1}t_{x}T_{0}U_{0}=S_{x}^{-1}T_{x}U_{0}$ is a non-trivial coset of $S_{0}T_{0}U_{0}$ in $H$, a contradiction.
So, if $|S||T||U|>|G|=p|H|$ then $\sigma=\tau=\upsilon=p$ and $S,T,U\not\leq H$ and $|S_{0}T_{0}U_{0}|<|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{3}}<|H|$. Also, as $H$ is maximal in $G$ and $U\not\leq H$ it follows that $U_{0}=U\cap H\trianglelefteq\langle U,H\rangle=G$ (and in the same way $T_{0}=T\cap H\trianglelefteq\langle T,H\rangle=G$). Now applying Lemma 3.3 shows that $\{S_{x}^{-1}T_{x}U_{0}\}_{xH\in G/H}$ and $\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H}$ are each collections of $p$ cosets of $S_{0}T_{0}U_{0}$ in $H$ having only the trivial coset $S_{0}T_{0}U_{0}$ in common. This means that $S_{0}T_{0}U_{0}$ has at least (2p1)(2p-1) cosets in $H$, i.e. $|H:S_{0}T_{0}U_{0}|=\frac{|H|}{|S_{0}||T_{0}||U_{0}|}\geq 2p-1$, which means that $|S||T||U|=p^{3}|S_{0}||T_{0}||U_{0}|\leq p^{3}\cdot\frac{|H|}{(2p-1)}=p^{3}\cdot\frac{1}{p(2p-1)}|G|$$\frac{p^{2}}{2p-1}|G|$.
A simple corollary is noted for some special cases,
@@
If $G$ is a group with an abelian normal subgroup of prime index $p$ such that $(2p-1)\nmid|G|$, that is, (2p1)(2p-1) does not divide $|G|$, then $\rho_{0}(G)\leq\frac{1}{2}p$. If $G$ is a $p$-group with an abelian subgroup of index $p$ then $\rho_{0}(G)=1$.
@@
In the first case, if $H\leq G$ is abelian and normal of prime index $p$ such that $(2p-1)\nmid|G|$ then $(2p-1)\nmid|H|$, and if $(S,T,U)$ is a subgroup TPP triple of $G$ then the subgroup $S_{0}T_{0}U_{0}<H$, as defined above, cannot be as large as $\frac{|H|}{(2p-1)}$, i.e. $|S_{0}||T_{0}||U_{0}|<\frac{|H|}{(2p-1)}$, which means that $|S||T||U|=p^{3}|S_{0}||T_{0}||U_{0}|\leq p^{3}\cdot\frac{|H|}{2p}=\frac{1}{2}p|G|$. In the second case, if $G$ is a $p$-group then the product of the orders of any of its subgroups is a $p$-power, and if $G$ has an abelian subgroup of index $p$ (which is thus necessarily normal in $G$), and $(S,T,U)$ is a subgroup TPP triple of $G$, then $|S||T||U|$ must be a $p$-power such that $|S||T||U|<\frac{p^{2}}{2p-1}|G|<p|G|$, i.e. $|S||T||U|\leq|G|$. ∎
The contrapositive of this corollary explains the non-trivial $\rho_{0}$ values computed by Hedtke and Murthy for all the groups listed in [5, Tables 1-4] which satisfy the conditions of the corollary while achieving $\rho_{0}>1$ or $\rho_{0}>\frac{1}{2}p$, and this has been verified by GAP computations in these groups.
5. Analysis and Conclusion

### v7

@@ 3. Coset Decomposition of TPP Triples
(2) Let $xH,yH\in\overline{S}\cap\overline{T}$ be distinct, and let $U_{0}\trianglelefteq G$. As in part (1) elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$, $s_{y}\in S_{y}$ and $t_{y}\in T_{y}$ can be chosen to form cosets $s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})=S_{x}^{-1}T_{x}U_{0}$ and $s_{y}^{-1}t_{y}(S_{0}T_{0}U_{0})=S_{y}^{-1}T_{y}U_{0}$ of $S_{0}T_{0}U_{0}$ in $H$, but these are distinct by the distinctness of $xH,yH$. To see this, suppose there are elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$, $s_{y}\in S_{y}$, $t_{y}\in T_{y}$, $u_{0},\widetilde{u}_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}t_{y}\widetilde{u}_{0}$. This can be rearranged as $s_{x}s_{y}^{-1}t_{y}\widetilde{u}_{0}u_{0}^{-1}t_{x}^{-1}=1$. As $U_{0}$ is normal in $G$ there is a $\hat{u_{0}}\in U_{0}$ such that $\widetilde{u}_{0}u_{0}^{-1}t_{x}^{-1}=t_{x}^{-1}\hat{u}_{0}$. Substituting this into the equation above implies that $s_{x}s_{y}^{-1}t_{y}t_{x}^{-1}\hat{u}_{0}=1$. However the TPP for $(S,T,U)$ implies that $s_{x}=s_{y}$, $t_{x}=t_{y}$, a contradiction.
Note here that it is necessary to assume the cosets $xH,yH$ are distinct, and that $U_{0}$ is normal in $G$, unlike in other parts of the proof. Also note that by transposing $T$ and $U$ in the argument above, it can be shown that cosets of the form $S_{x}^{-1}U_{x}T_{0}$ and $S_{y}^{-1}U_{y}T_{0}$ are distinct if $xH,yH\in\overline{S}\cap\overline{U}$ are distinct, and $T_{0}$ is normal in $G$.
(3) Suppose $xH\in\overline{S}\cap\overline{T}$ and $yH\in\overline{S}\cap\overline{U}$, and that $xH$ is non-trivial, that is, $xH\neq H$. As above there are cosets $S_{x}^{-1}T_{x}U_{0}$ and $S_{y}^{-1}U_{y}T_{0}$ of $S_{0}T_{0}U_{0}$ in $H$, but these are distinct. To see this, suppose there are elements $s_{x}\in S_{x}$, $s_{y}\in S_{y}$, $t_{x}\in T_{x}$, $u_{y}\in U_{y}$, $t_{0}\in T_{0}$, and $u_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}u_{y}t_{0}$, or $s_{x}s_{y}^{-1}u_{y}t_{0}u_{0}^{-1}t_{x}^{-1}=1$. Since $t_{0}$ and $u_{0}^{-1}$ commute in $H$ this can be rewritten as $s_{x}s_{y}^{-1}u_{y}u_{0}^{-1}t_{0}t_{x}^{-1}=1$. However the TPP for $S,U,T$ implies that $t_{x}=t_{0}$, a contradiction if $xH\neq H$. In the same way, the assumption that $yH$ is non-trivial would also lead to a similar contradiction. Note here that there is no need to assume distinctness of the cosets $xH,yH$, only that at least one of them is non-trivial.
4. The Main Result for Subgroup TPP Ratio
@@
If $G$ is a group with an abelian normal subgroup $H$ of prime index $p$ then
(4.1) $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$
where equality $\rho_{0}(G)=\frac{p^{2}}{2p-1}$ implies that 2p12p-1 divides $|H|$, and |G|=p(2p1)m|G|=p(2p-1)m where $m$ is the order of the proper subgroup of $H$ associated with the $H$-restriction of the subgroup TPP triple of $G$ through which the equality is achieved.
@@
Let $G$ and the prime $p$ be as given. Let $H$ be an abelian normal subgroup of index $p$ (thus making it maximal in $G$). Then $G/H$ is a cyclic group of order $p=|G:H|=\frac{|G|}{|H|}$ whose elements are $p$ (left) cosets $xH$, for $x\in G$. Let $(S,T,U)$ be a subgroup TPP triple of $G$ with $H$-supports $\overline{S},\overline{T},\overline{U}$ as defined in Observation 3.1. By part (2) of that Observation $\overline{S},\overline{T},\overline{U}\leq G/H$, and letting $\sigma=|\overline{S}|$, $\tau=|\overline{T}|$, $\upsilon=|\overline{U}|$ then $\sigma,\tau,\upsilon\in\{1,p\}$. The sets $S_{x},T_{y},U_{z}$, as defined in Lemma 3.3 will also be used at certain points: these are defined as $S_{x}:=S\cap xH$, $T_{y}:=T\cap yH$, $U_{z}:=U\cap zH$ for any $xH,yH,zH\in G/H$. By definition, if $xH\in G/H$ then $xH\in\overline{S}\iff S_{x}\neq\emptyset$, and there are analogues for $T_{y}$ and $U_{z}$.
@@
Let $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$. Then $(S_{0},T_{0},U_{0})$ is a subgroup TPP triple of $H$ (by Observation 2.4) of size $|S_{0}||T_{0}||U_{0}|=|S_{0}T_{0}U_{0}|=\frac{|S||T||U|}{\sigma\tau\upsilon}\leq|H|$ and $|S||T||U|\leq\frac{\sigma\tau\upsilon}{p}|G|$ (by Proposition 2.5 and Observation 3.2). Note that $S_{0}T_{0}U_{0}\trianglelefteq H$ (as $H$ is abelian), and by the Second Isomorphism Theorem, $S_{0}\trianglelefteq S$, $T_{0}\trianglelefteq T$, and $U_{0}\trianglelefteq U$.
Observation 3.2 implies the general bound $|S||T||U|\leq\frac{\sigma\tau\upsilon}{p}|G|$ applies, independently of any special assumptions about $|S||T||U|$. If at least two of $\sigma,\tau,\upsilon$ are equal to $1$ then $|S||T||U|\leq\frac{\sigma\tau\upsilon}{p}|G|\leq|G|$, and there is nothing to prove. So to achieve $|S||T||U|>|G|$ there are only two cases to consider: (i) $\sigma\tau\upsilon=p^{2}$ or (ii) $\sigma\tau\upsilon=p^{3}$.
(*) Now the special assumption is introduced (for which a contradiction is sought): suppose $|S||T||U|>\frac{p^{2}}{2p-1}|G|=\frac{p^{3}}{2p-1}|H|$.
Consider case (i): let $\sigma=\tau=p$ and $\upsilon=1$, without loss of generality (permutation invariance of $S,T,U$ by Observation 2.3), which means that $S$ and $T$ are supported on all $p$ cosets of $H$, and $U$ only on $H$ ($U=U_{0}$). Then $|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{2}}>\frac{p}{2p-1}|H|$ and $\frac{|H|}{|S_{0}||T_{0}||U_{0}|}=|H:S_{0}T_{0}U_{0}|<\frac{2p-1}{p}=2-\frac{1}{p}<2$, that is, $|H:S_{0}T_{0}U_{0}|=1$ and $S_{0}T_{0}U_{0}=H$. However, from the p1p-1 non-trivial cosets of $H$ on which $S$ and $T$ are commonly supported any coset $xH$ and elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$ can be chosen to form a coset $s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})=S_{x}^{-1}T_{x}U_{0}$ of $S_{0}T_{0}U_{0}$ in $H$. And, by Lemma 3.3 part (1) this is non-trivial, that is, $S_{x}^{-1}T_{x}U_{0}\neq S_{0}T_{0}U_{0}$, a contradiction. This argument also shows that when only two of the subgroups $S,T,U$ are supported on all $p$ cosets of $H$ then $|S||T||U|<\frac{p^{2}}{2p-1}|G|$ and $\rho_{0}(G)$ can never achieve $\frac{p^{2}}{2p-1}$.
Consider case (ii), that is, $\sigma=\tau=\upsilon=p$, meaning that $S,T,U$ are supported on all $p$ cosets of $H$, or, equivalently, $\overline{S}=\overline{T}=\overline{U}=G/H$ and the sets $S_{x},T_{y},U_{z}$, as defined above, are all non-empty. Importantly, here, the normality of $T_{0}$ and $U_{0}$ in $G$ is necessary, and, indeed, follows from the conditions here: to see this, note that as $T_{0}$ is normal in both $T$ and $H$ it is normal in the subgroup $\langle T,H\rangle$ generated by them, and as $T$ is not contained in $H$, where $H$ is maximal in $G$, then $\langle T,H\rangle=G$. In the same way, $U_{0}$ is normal in $G$.
Now, $|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{3}}>\frac{1}{2p-1}|H|$, that is, $|H:S_{0}T_{0}U_{0}|<2p-1$ and $S_{0}T_{0}U_{0}$ has less than 2p12p-1 cosets in $H$. However, using Lemma 3.3 parts (2)-(3) two collections of p1p-1 non-trivial cosets each of $S_{0}T_{0}U_{0}$ in $H$ can be formed, namely, $\mathcal{A}=\{S_{x}^{-1}T_{x}U_{0}\}_{xH\in G/H\backslash\{H\}}$ and $\mathcal{B}=\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H\backslash\{H\}}$, that are mutually distinct, that is, $S_{x}^{-1}T_{x}U_{0}\neq S_{y}^{-1}U_{y}T_{0}$ for any pair $S_{x}^{-1}T_{x}U_{0}\in\mathcal{A}$, $S_{y}^{-1}U_{y}T_{0}\in\mathcal{B}$. And this implies that the total number of cosets of $S_{0}T_{0}U_{0}$ in $H$, including the trivial coset $S_{0}T_{0}U_{0}$, is equal to 1+2(p1)=2p11+2(p-1)=2p-1, a contradiction.
Thus (*) cannot hold and so $|S||T||U|\leq\frac{p^{2}}{2p-1}|G|$, as claimed.
For the final part of the theorem, suppose $|S||T||U|=\frac{p^{2}}{2p-1}|G|$. Then $|S_{0}||T_{0}||U_{0}|=\frac{1}{2p-1}|H|$ and $|H:S_{0}T_{0}U_{0}|=2p-1$ (so $2p-1\mathrel{|}|H|$ also). As $S_{0}T_{0}U_{0}\leq G$ then $|G:S_{0}T_{0}U_{0}|=|G:H||H:S_{0}T_{0}U_{0}|=p(2p-1)$ and $|G|=p(2p-1)|S_{0}T_{0}U_{0}|=p(2p-1)|S_{0}||T_{0}||U_{0}|$. ∎
Note that the upper bound $\frac{p^{2}}{2p-1}$ here for $\rho_{0}$ is an improvement by a factor of $\frac{1}{2p-1}$ on the general upper bound of $p^{2}$ noted previously which holds when $G$ contains any abelian subgroup of index $p$ (note that $\frac{p^{2}}{2p-1}\longrightarrow\left(\frac{1}{2}p\right)^{+}$ as $p\longrightarrow\infty$).
Example 4.2.
A simple illustration of the theorem (and the idea of coset decomposition) is given below using the dihedral group $D_{2n}$ of order $2n$ where $n$ is divisible by $3$, that is, $n=3m$ for some positive integer $m$ (so that $|D_{2n}|=6m$), and the equality $\rho_{0}(G)=\frac{p^{2}}{2p-1}$ is achieved for $p=2$.
Using the standard presentation $\langle r,f\mathrel{|}r^{n}=r^{3m}=1,f^{2}=1,frf=r^{-1}\rangle$ of $D_{2n}$, and letting $H$ be the cyclic subgroup $\langle r\rangle=\{r^{0}=1,r,\ldots,r^{3m-1}\}$ of order $n=3m$ (and index $2$), and $fH=\{f,fr,\ldots,fr^{3m-1}\}$ its non-trivial coset, consider the subgroups
$S$ $=\langle r^{3},f\rangle=\{r^{3i}f^{j}\mathrel{|}i\in\mathbb{Z}_{m},j\in\mathbb{Z}_{2}\}$
$T$ $=\langle fr\rangle=\{1,fr\}$
$U$ $=\langle fr^{2}\rangle=\{1,fr^{2}\}$
These subgroups are generated by the set $\{r^{3},f\}$ and the involutions $fr$ and $fr^{2}$ respectively. The largest of these, $S$, is generated by two elements, one, $r^{3}$, of order $\frac{1}{3}n=m$, and the other, $f$, an involution, and so forms of a subgroup of order $\frac{2}{3}n=2m$. Clearly, $T\cap U=\{1\}$, and it is easy to check that $TU=\{1,fr,fr^{2},frfr^{2}=r^{-1}r^{2}\}=\{1,fr,fr^{2},r\}$ and $S\cap TU=\{1\}$. Thus $(S,T,U)$ is a (subgroup) TPP triple of $D_{2n}$ of type $(\frac{2}{3}n,2,2)=(2m,2,2)$ and size $\frac{8}{3}n=8m$, thus achieving $\rho_{0}(D_{2n})=\frac{4}{3}$.
[…]
$|S||T||U|$ $=(|S_{0}|+|S_{1}|)(|T_{0}|+|T_{1}|)(|U_{0}|+|U_{1}|)$
$=\sum_{0\leq i,j,k\leq 1}|S_{i}||T_{j}||U_{k}|$
$=\sum_{0\leq i,j,k\leq 1}|s_{i}^{-1}t_{j}u_{k}(S_{0}T_{0}U_{0})|$
$=\sum_{0\leq i,j,k\leq 1}|S_{0}||T_{0}||U_{0}|$
$=\sum_{0\leq i,j,k\leq 1}|S_{0}|$
$=8m=\frac{8}{3}n=\frac{4}{3}|D_{2n}|$
where $|S_{i}||T_{j}||U_{k}|=|S_{i}|=|S_{0}|$ is the size of the TPP triple $(S_{i},T_{j},U_{k})$, and elements $s_{i}\in S_{i}$, $t_{j}\ \in T_{j}$, $u_{k}\in U_{k}$ can be chosen appropriately for $i,j,k\in\{0,1\}$. There are thus $8$ of these, each contributing $m$ to the size of $|S||T||U|$.
A simple corollary is noted for some special cases.
@@
(1) If $G$ is a group with an abelian normal subgroup $H$ of prime index $p$ and $(2p-1)\nmid|G|$, that is, (2p1)(2p-1) does not divide $|G|$, then $\rho_{0}(G)\leq\frac{1}{2}p$. (2) If $G$ is a $p$-group with an abelian subgroup of index $p$ then $\rho_{0}(G)=1$.
@@
(1) By Theorem 4.1, $|S||T||U|\leq\frac{p^{2}}{2p-1}|G|=\frac{p^{3}}{2p-1}|H|$ holds for any subgroup TPP triple $(S,T,U)$ of $G$, where $G$ and $H$ are given as above. If $(2p-1)\nmid|G|$ then $(2p-1)\nmid|H|$ and $|S||T||U|<\frac{p^{2}}{2p-1}|G|$, which implies that $|S||T||U|\leq\frac{p^{2}}{2p}|G|=\frac{1}{2}p|G|$. This shows that if 2p12p-1 does not divide $|G|$ then $\rho_{0}(G)=1$ when $p=2$, and $\rho_{0}(G)>1$ is possible only if $p\geq 3$.
(2) Let $G$ be a (finite) $p$-group with an abelian subgroup $H$ of index $p$. Then $H$ is maximal in $G$, and thus normal [1, Proposition 3, p.73]. And the size of any subgroup TPP triple of $G$ is a $p$-power. By Theorem 4.1 if $(S,T,U)$ is a subgroup TPP triple of $G$, then $|S||T||U|$ must be a $p$-power such that $|S||T||U|<\frac{p^{2}}{2p-1}|G|<p|G|$, that is, $|S||T||U|\leq|G|$. ∎
The contrapositive of this corollary explains the non-trivial $\rho_{0}$ values computed by Hedtke and Murthy for all the groups listed in [5, Tables 1-4] that achieve $\rho_{0}>1$ or $\rho_{0}>\frac{1}{2}p$, and this has been verified by GAP computations in these groups. It also explains the trivial $\rho_{0}$ values listed in those tables for the $p=2$ case where either $2\cdot 2-1=3$ does not divide the order of an abelian subgroup of index $2$, as in the case of the dihedral groups $D_{2n}$ where $3$ does not divide $n$, or the group is a $2$-group containing an abelian subgroup of index $2$.
5. Concluding Remarks